# HERGC: Heterogeneous Experts Representation and Generative Completion for Multimodal Knowledge Graphs

### v1

@@ Abstract
Multimodal knowledge graphs (MMKGs) enrich traditional knowledge graphs (KGs) by incorporating diverse modalities such as images and text. multimodal knowledge graph completion (MMKGC) seeks to exploit these heterogeneous signals to infer missing facts, thereby mitigating the intrinsic incompleteness of MMKGs. Existing MMKGC methods typically leverage only the information contained in the MMKGs under the closed-world assumption and adopt discriminative training objectives, which limits their reasoning capacity during completion. Recent generative completion approaches powered by advanced large language models (LLMs) have shown strong reasoning abilities in unimodal knowledge graph completion, but their potential in MMKGC remains largely unexplored. To bridge this gap, we propose HERGC, a Heterogeneous Experts Representation and Generative Completion framework for MMKGs. HERGC first deploys a Heterogeneous Experts Representation Retriever that enriches and fuses multimodal information and retrieves a compact candidate set for each incomplete triple. It then uses a Generative LLM Predictor fine-tuned on minimal instruction data to accurately identify the correct answer from these candidates. Extensive experiments on three standard MMKG benchmarks demonstrate HERGC’s effectiveness and robustness, achieving state-of-the-art performance.
@@ 1 Introduction
Like traditional KGs, MMKGs also suffer from incompleteness, often due to missing facts in the underlying data sources or facts that have yet to be discovered by humans. Unlike conventional knowledge graph completion (KGC), which primarily leverages the graph’s topological structure and local neighborhood information, multimodal knowledge graph completion (MMKGC) (Chen et al., 2024b) introduces additional complexity through the incorporation of multimodal signals. MMKGC has advanced considerably in recent years, with most work concentrating on modality fusion (Li et al., 2023; Shang et al., 2024) and modality information representation (Zhang et al., 2025a). These approaches ultimately produce yield joint triple embeddings or employ ensemble strategies to score candidate facts. While these MMKGC methods offer valuable insights, their reasoning capabilities during the completion process remain limited due to their reliance solely on closed-world triples and available multimodal information.
@@
Meanwhile, recent advances in unimodal KGC have introduced generative completion approaches powered by large language models (LLMs) Wei et al. (2023); Zhang et al. (2024); Liu et al. (2024). By fine-tuning or leveraging in-context learning, these approaches exploit the extensive factual knowledge and reasoning capabilities that LLMs acquire during pre-training, achieving state-of-the-art results in KGC. However, this generative paradigm remains largely unexplored within the MMKGC setting. In particular, designing mechanisms to integrate multimodal information to guide LLM reasoning and constrain the search space of its responses remains a critical open challenge.
@@
To address these challenges and fill the gap in generative MMKGC, we propose HERGC, a novel generative framework that overcomes the limitations of existing MMKGC approaches that rely on MMKG’s own contents for reasoning. Figure 1 briefly illustrates the differences between our HERGC and existing methods. Drawing on the retrieval-augmented generation (RAG) idea (Lewis et al., 2020), HERGC comprises two core component: the Heterogeneous Experts Representation Retriever (HERR) and the Generative LLM Predictor (GLP). HERR employs a Mixture of Heterogeneous Experts Network (MoHE) to enrich each modality’s embeddings from multiple and hierarchical perspectives and a Relation-aware Gated Multimodal Unit (RaGMU) to obtain high-quality fused embeddings, which are then used to score and retrieve candidate entities. GLP injects the fused multimodal embeddings into the LLM and performs LoRA fine-tuning on minimal instruction data, enabling the LLM to accurately select the correct entity from the retrieved candidates. We conduct comprehensive experiments on three public MMKG benchmarks to validate the effectiveness and robustness of HERGC. Our contributions are summarized as follows:
  • We propose HERGC, the first, to the best of our knowledge, MMKGC framework based on the generative paradigm.
  • We developed a novel ranking-based retrieval module, HERR, which combines the Mixture of Heterogeneous Experts Network and the Relation-aware Gated Multimodal Unit to effectively extract cross-modal information and produce high-quality fused multimodal embeddings.
@@
  • We perform extensive experiments on three standard MMKGC benchmarks, showing that HERGC consistently outperforms strong baselines across all datasets.
@@ 2 Related Work
2.1 Tradition Knowledge Graph Completion
Tradition Knowledge Graph Completion (KGC) primarily focuses on embedding entities and relations into continuous vector spaces to predict triples. Most of them leverage KG’s structure and design various score function to learn the embedding by maximize the positive and negative samples score difference, such as Translational-Distance approaches (TransE (Bordes et al., 2013) and RotatE (Sun et al., 2019)) and Semantic-Matching approaches (DistMult (Yang et al., 2014), ComplEX Trouillon et al. (2016), and Tucker Balažević et al. (2019)).To further consider neighbor aggregation and message passing to improve the representation power of embedding, graph neural network based (GNN-based) methods have been proposed, such as R-GCN (Schlichtkrull et al., 2018) and CompGCN (Vashishth et al., 2019). Besides structural information, KG, as a semantic network, naturally carries text information. Therefore, the text-based method that mainly uses text information, which encodes text information in KG through a pretrained language model (PLM), has been proposed, including KG-Bert (Yao et al., 2019) and SimKGC (Wang et al., 2022). With the recent development of LLM, the novel generative methods have come into view. They mainly use the rich external knowledge and powerful reasoning capabilities of LLMs to complete KGC in a sequence-to-sequence form, including KICGPT (Wei et al., 2023), KoPA (Zhang et al., 2024) and DIFT (Liu et al., 2019).
@@ 3 Methodology
3.1 Preliminary
Multimodal Knowledge graph (MMKG). In this work, we focus on the most common form of MMKGs with dual visual and textual modalities. An MMKG can be represented as a directed multigraph with modal attributes, denoted as $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T},\mathcal{V},\mathcal{D})$, where $\mathcal{E}$ is the set of entities, $\mathcal{R}$ is the set of relations, and $\mathcal{T}=\{(h,r,t)|h,t\in\mathcal{E},t\in\mathcal{R}\}$ is the set of triples (i.e. (head entity, relation, tail entity)). The $\mathcal{V}$ and $\mathcal{D}$ are the collection of visual images and descriptive text associated with entities.
Multimodal Knowledge graph Completion (MMKGC). MMKGC aims to make full use of the observed triples $\mathcal{T}$ together with the visual and textual attributes of entities ($\mathcal{V}$ and $\mathcal{D}$) to infer missing triples. The set of potential facts is defined as $\{(h^{\prime},r^{\prime},t^{\prime})\mid h^{\prime},t^{\prime}\in\mathcal{E},\ r^{\prime}\in\mathcal{R}\}$, where $(h^{\prime},r^{\prime},t^{\prime})\notin\mathcal{T}$ represents missing triples in the MMKG. In this work, we formulate MMKGC as the task of completing incomplete query triples of the form (?$,r_{q},t_{q})$ and $(h_{q},r_{q},?)$, corresponding to head prediction and tail prediction, respectively. Here, we refer $h_{q}$ or $t_{q}$ the query entity and $r_{q}$ the the query relation.
@@ 3.2 Multimodal Information Embedding
To enable effective fusion of multimodal information, we first encode each entity’s visual and textual modalities into embedding representations, denoted as $\mathbf{e}_{\mathcal{V}}$ and $\mathbf{e}_{\mathcal{D}}$, respectively. In addition to these modalities, the graph structure in the KG provides valuable graph contextual cues. Therefore, we also embed the structural information of each entity, resulting in a representation $\mathbf{e}_{\mathcal{S}}$.
Image and Text Embedding. We utilize pre-trained models to encode the visual and textual information associated with each entity. Specifically, we adopt BERT Devlin et al. (2019), an encoder-only transformer model trained on large-scale textual corpora, to obtain text embeddings, and VGG (Simonyan and Zisserman, 2014), a convolutional neural network trained on large-scale image datasets, to extract visual features. Each entity’s descriptive text and image are processed through BERT and VGG, respectively, yielding the initial modality-specific embeddings $\mathbf{e}_{\mathcal{D}}$ and $\mathbf{e}_{\mathcal{V}}$.
Structure Embedding. To encode structural information from the MMKGC, we adopt TuckER (Balažević et al., 2019), a representative KGE model that learns entity and relation embeddings based on tensor factorization. We select TuckER to ensure consistency with the scoring function used in our retrieval module. The resulting embedding is taken as the structural representation $\mathbf{e}_{\mathcal{S}}$ for each entity.
@@ 3.3 Heterogeneous Experts Representation Retriever
The retriever in KGC is typically a representation model that scores and ranks candidate triples. Under the complex multimodal setting, however, maintaining high-quality entity representations and retrieval sets becomes challenging. To address this, we introduce the Heterogeneous Experts Representation Retriever (HERR). HERR first enhances modality-specific features using a Mixture of Heterogeneous Experts network. It then fuses these features via a Relation-aware Gated Multimodal Unit to produce fused embeddings. Finally, HERR employs a scoring function to compute triple scores and generate a ranked candidate list.
Mixture of Heterogeneous Experts Module. To obtain representative and multi-perspective embeddings for each modality we design a heterogeneous experts layer composed of both simple and complex experts. Given the inherent heterogeneity of modality-specific embeddings, our goal is to align them across modalities while preserving their unique characteristics during fusion. To this end, we introduce the Mixture of Heterogeneous Experts (MoHE) module.
The standard Mixture of Experts (MoE) (Shazeer et al., 2017) architecture consists of a set of independent expert networks and a gating network that determines expert assignments. Specifically, given the input feature vector of the modality, MoE outputs the weighted sum of top-$\kappa$ experts outputs:
@@
The simplest expert in the MoE layer performs only a linear transformation and feature whitening. To better adapte heterogeneous modalities, we augment the expert set in MoHE with complex PHM experts, which are based on Block-Hypercomplex Linear Transformations (Zhang et al., 2021). Specifically, the input $\mathbf{x}\in\mathbb{R}^{d}$ will be partitioned into $n$ sub-blocks of size $d/n$:
@@
Relation-aware Gated Multimodal Units.
Gated Multimodal Units (Arevalo et al., 2017) provide a "multimodal projection $\to$ gating $\to$ weighted fusion" idea to combine modality-specific representations. To tailor this mechanism to MMKGC, we introduced the Relation-aware Gated Multimodal Units, denoted RaGMU.
@@
$\mathbf{h}_{m}=\tanh(\mathbf{W}_{proj,m}\mathbf{x}+\mathbf{b}_{proj,m})$ (6)
where $\mathbf{W}_{proj,m}$ and $\mathbf{b}_{proj,m}$ are the project matrix and basis of RaGMU projector for modality $m$.
@@
$\mathbf{z}=\mathrm{softmax}{(r_{z}({\mathbf{W}_{z}\mathbf{h}_{concat}+\mathbf{b}_{z}}))}$ (7)
where $\mathbf{h}_{concat}=[\mathbf{h}_{1};\mathbf{h}_{2};...;\mathbf{h}_{M}]$ is the concatenation of all projected modality embeddings, the $\mathbf{W}_{z}$ and $\mathbf{b}_{z}$ are the gating weight matrix and bias, and the $r_{z}$ is the scalar that injects relation awareness.
Finally, the fused multimodal embedding is obtained by applying the gate vector to each modality’s hidden projection $\mathbf{h}_{m}$:
@@
where $\odot$ denotes Hadamard product and $\mathbf{z}_{m}$ is the $m$-th element gate vector corresponding to modality $m$.
Score Function. After getting the fused multimodal embeddings, we compute triple plausibility scores using the TuckER scoring function:
@@
where $\mathbf{h}_{h}$ and $\mathbf{h}_{t}$ denote the fused embeddings of the head $h$ and tail $t$, $\mathbf{r}_{r}$ denotes the embedding of the relation $r$, and ×$\times_{n}$ denotes the $n$-mode tensor product (Balažević et al., 2019).
To train the model, we adopt a binary classification objective that encourages higher scores for positive triples and lower scores for negative ones. The loss function is defined as:

### v2

@@ Abstract
Multimodal knowledge graphs (MMKGs) enrich traditional knowledge graphs (KGs) by incorporating diverse modalities such as images and text. multimodal knowledge graph completion (MMKGC) seeks to exploit these heterogeneous signals to infer missing facts, thereby mitigating the intrinsic incompleteness of MMKGs. Existing MMKGC methods typically leverage only the information contained in the MMKGs under the closed-world assumption and adopt discriminative training objectives, which limits their reasoning capacity during completion. Recent large language models (LLMs), empowered by massive parameter scales and pretraining on vast corpora, have demonstrated strong reasoning abilities across various tasks. However, their potential in MMKGC remains largely unexplored. To bridge this gap, we propose HERGC, a flexible Heterogeneous Experts Representation and Generative Completion framework for MMKGs. HERGC first deploys a Heterogeneous Experts Representation Retriever that enriches and fuses multimodal information and retrieves a compact candidate set for each incomplete triple. It then uses a Generative LLM Predictor, implemented via either in-context learning or lightweight fine-tuning, to accurately identify the correct answer from these candidates. Extensive experiments on three standard MMKG benchmarks demonstrate HERGC’s effectiveness and robustness, achieving superior performance over existing methods.
@@ 1 Introduction
Like traditional KGs, MMKGs also suffer from incompleteness, often due to missing facts in the underlying data sources or facts that have yet to be discovered by humans. Unlike unimodal knowledge graph completion (KGC), which primarily leverages the graph’s topological structure and local neighborhood information, multimodal knowledge graph completion (MMKGC) (Chen et al., 2024b) introduces additional complexity through the incorporation of multimodal signals. MMKGC has advanced considerably in recent years, with most work concentrating on modality fusion (Li et al., 2023; Shang et al., 2024) and modality information representation (Zhang et al., 2025a). These approaches ultimately yield joint triple embeddings or employ ensemble strategies to score candidate facts. While these MMKGC methods offer valuable insights, their reasoning capabilities during the completion process remain limited due to their reliance solely on closed-world triples and available multimodal information.
@@
Meanwhile, recent advances in unimodal KGC have introduced generative completion approaches powered by large language models (LLMs) Wei et al. (2023); Zhang et al. (2024); Liu et al. (2024). By leveraging in-context learning or fine-tuning, these approaches exploit the extensive factual knowledge and reasoning capabilities that LLMs acquire during pre-training, achieving strong performance. However, this generative paradigm remains largely underexplored in the MMKGC setting due to its inherent limitations. First, these methods typically rely on existing knowledge graph embedding (KGE) models (e.g., RotatE Sun et al. (2019), as used in KICGPT Wei et al. (2023)) for candidate retrieval, which ignore multimodal signals and yield low-recall candidate sets. Second, their LLM-based predictors are designed to process only textual inputs, excluding visual and structural modalities that are essential for comprehensive multimodal reasoning. These limitations motivate the development of novel mechanisms that seamlessly integrate multimodal information into both retrieval and generation, enabling more accurate, context-aware reasoning within constrained candidate spaces.
@@
To address these challenges and fill the gap in generative MMKGC, we propose HERGC, a novel and flexible generative framework that overcomes the limitations of closed-world reasoning based solely on MMKG contents and introduces a multimodal-aware generative completion paradigm. Figure 1 briefly illustrates the differences between our HERGC and existing MMKGC methods. Inspired by the retrieval-augmented generation (RAG) idea (Lewis et al., 2020), HERGC comprises two core component: the Heterogeneous Experts Representation Retriever (HERR) and the Generative LLM Predictor (GLP). HERR employs a Mixture of Heterogeneous Experts Network (MoHE) to enrich each modality’s embeddings from multiple and hierarchical perspectives and a Relation-aware Gated Multimodal Unit (RaGMU) to obtain high-quality fused embeddings, which are then used to score and retrieve candidate entities. GLP supports $(1)$ directly using powerful closed-source LLMs (e.g., GPT-4) via APIs to make predictions with in-context learning, or $(2)$ injecting the fused multimodal embeddings into open-source LLMs (e.g., LLaMA) and performing LoRA fine-tuning on minimal instruction data. This flexible design allows GLP to adapt to diverse resource conditions while enabling the LLMs to accurately select the correct entity from the retrieved candidates. We conduct comprehensive experiments on three public MMKG benchmarks to validate the effectiveness and robustness of HERGC. Our contributions are summarized as follows:
  • We propose HERGC, the first, to the best of our knowledge, MMKGC framework based on the generative paradigm. It features a flexible Generative LLM Predictor (GLP) that supports both open-source and closed-source LLMs, enabling effective integration of external knowledge for complex multimodal structural reasoning.
  • We design a novel Heterogeneous Experts Representation Retriever (HERR), which combines a Mixture of Heterogeneous Experts (MoHE) and a Relation-Aware Gated Multimodal Unit (RaGMU) to extract multi-perspective signals from heterogeneous and distributionally diverse modalities, and to adaptively fuse them based on relation types, producing high-quality fused embeddings and candidate sets.
@@
  • We conduct extensive experiments on three standard MMKGC benchmarks, demonstrating that HERGC consistently outperforms strong baselines and exhibits robust performance across diverse settings.
@@ 2 Related Work
Unimodal Knowledge Graph Completion (KGC) primarily focuses on embedding entities and relations into continuous vector spaces to predict triples. Most of them leverage KG’s structure and design various score function to learn the embedding by maximize the positive and negative samples score difference, such as Translational-Distance approaches (TransE (Bordes et al., 2013) and RotatE (Sun et al., 2019)) and Semantic-Matching approaches (DistMult (Yang et al., 2015), ComplEX Trouillon et al. (2016), and Tucker Balažević et al. (2019)).To improve the representation power of embedding, graph neural network based (GNN-based) methods have been proposed, such as R-GCN (Schlichtkrull et al., 2018) and CompGCN (Vashishth et al., 2020). Besides structural information, KG, as a semantic network, naturally carries text information. Therefore, the text-based method that mainly uses text information, which encodes text information in KG through a pretrained language model (PLM), has been proposed, including KG-Bert (Yao et al., 2019) and SimKGC (Wang et al., 2022). With the recent development of LLM, the novel generative methods have come into view. They mainly use the rich external knowledge and powerful reasoning capabilities of LLMs to complete KGC in a sequence-to-sequence form, including KICGPT (Wei et al., 2023), KoPA (Zhang et al., 2024) and DIFT (Liu et al., 2019).
@@ 2.1 Multimodal Knowledge Graph Completion
3 Preliminary
In this work, we focus on the most common form of Multimodal Knowledge graph (MMKG) with dual visual and textual modalities. An MMKG can be represented as a directed multigraph with modal attributes, denoted as $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T},\mathcal{V},\mathcal{D})$, where $\mathcal{E}$ is the set of entities, $\mathcal{R}$ is the set of relations, and $\mathcal{T}=\{(h,r,t)|h,t\in\mathcal{E},t\in\mathcal{R}\}$ is the set of triples (i.e. (head entity, relation, tail entity)). The $\mathcal{V}$ and $\mathcal{D}$ are the collection of visual images and descriptive text associated with entities.
Multimodal Knowledge graph Completion (MMKGC) aims to make full use of the observed triples $\mathcal{T}$ together with the visual and textual attributes of entities ($\mathcal{V}$ and $\mathcal{D}$) to infer missing triples. The set of potential facts is defined as $\{(h^{\prime},r^{\prime},t^{\prime})\mid h^{\prime},t^{\prime}\in\mathcal{E},\ r^{\prime}\in\mathcal{R}\}$, where $(h^{\prime},r^{\prime},t^{\prime})\notin\mathcal{T}$ represents missing triples in the MMKG. In this work, we formulate MMKGC as the task of completing incomplete query triples of the form $(?,r_{q},t_{q})$ and $(h_{q},r_{q},?)$, corresponding to head prediction and tail prediction, respectively. Here, we refer $h_{q}$ or $t_{q}$ the query entity and $r_{q}$ the the query relation.
@@ 4.1 Multimodal Information Embedding
To enable effective fusion of multimodal information, we first encode each entity’s visual and textual modalities into embedding representations, denoted as $\mathbf{e}_{\mathcal{V}}$ and $\mathbf{e}_{\mathcal{D}}$, respectively. Additionally, we embed each entity’s structural information from the KG into a structural representation $\mathbf{e}_{\mathcal{S}}$ to capture graph contextual cues.
Image and Text Embedding. We utilize pre-trained models to encode the visual and textual information associated with each entity. To ensure a fair comparison, we maintain consistency with recent baselines (Zhang et al., 2025b; Gao et al., 2025) by adopting BERT Devlin et al. (2019), an encoder-only transformer model trained on large-scale textual corpora, for text embeddings, and VGG (Simonyan and Zisserman, 2014), a convolutional neural network trained on large-scale image datasets, for visual feature extraction. Each entity’s descriptive text and image are processed through BERT and VGG, respectively, yielding the initial modality-specific embeddings $\mathbf{e}_{\mathcal{D}}$ and $\mathbf{e}_{\mathcal{V}}$.
Structure Embedding. To encode structural information from the MMKG, we adopt TuckER (Balažević et al., 2019), a representative KGE model that learns entity and relation embeddings via tensor factorization. TuckER is also employed as the scoring function in the retrieval module. The resulting embedding is taken as the structural representation $\mathbf{e}_{\mathcal{S}}$ for each entity.
@@ 4.2 Heterogeneous Experts Representation Retriever
The retriever in unimodal generative KGC typically relies on a simple KGE model to score and rank candidate triples. However, under the complex multimodal setting, maintaining high-quality entity representations and retrieval sets becomes significantly more challenging. To address this, we propose the Heterogeneous Experts Representation Retriever (HERR). HERR first enhances modality-specific features through a Mixture of Heterogeneous Experts network, then fuses them using a Relation-Aware Gated Multimodal Unit (RaGMU) to obtain joint embeddings. Finally, HERR employs a scoring function to compute triple scores and generate a ranked list of candidate entities.
Mixture of Heterogeneous Experts Module. To obtain representative and multi-perspective embeddings for each modality we design a heterogeneous experts layer composed of both simple and complex experts. Given the inherent heterogeneity and distinct distributions of modality-specific embeddings, our goal is to align them across modalities while preserving their unique characteristics during fusion. To this end, we introduce the Mixture of Heterogeneous Experts (MoHE) module.
MoHE extends the standard Mixture of Experts (MoE) (Shazeer et al., 2017) architecture by combining diverse expert types. Given the input feature vector of the modality, MoHE outputs the weighted sum of top-$\kappa$ experts outputs:
@@
The simple expert in the MoHE layer performs a linear transformation and feature whitening. To better adapt to heterogeneous modalities and capture richer cross-dimensional interactions, MoHE also incorporates complex PHM experts inspired by Block-Hypercomplex Linear Transformations (Zhang et al., ). Specifically, the input $\mathbf{x}\in\mathbb{R}^{d}$ will be partitioned into $n$ sub-blocks of size $d/n$:
@@
Relation-aware Gated Multimodal Units. In MMKGC tasks, the importance of each modality can vary across relation types. To address this, we propose the Relation-aware Gated Multimodal Unit (RaGMU), which dynamically adjusts fusion weights based on the relations.
@@
$\mathbf{h}_{m}=\tanh(\mathbf{W}_{proj,m}\mathbf{x}_{m}+\mathbf{b}_{proj,m})$ (6)
where $\mathbf{W}_{proj,m}$ and $\mathbf{b}_{proj,m}$ are the projection matrix and bias of RaGMU projector for modality $m$.
@@
$\mathbf{z}=\mathrm{softmax}{(g_{r}(\mathbf{r})\odot({\mathbf{W}_{z}\mathbf{h}_{concat}+\mathbf{b}_{z}}))}$ (7)
where $\mathbf{h}_{concat}=[\mathbf{h}_{1};\mathbf{h}_{2};...;\mathbf{h}_{M}]$ is the concatenation of all projected modality embeddings, the $\mathbf{W}_{z}$ and $\mathbf{b}_{z}$ are the gating weight matrix and bias, $\odot$ denotes Hadamard product, and $g_{r}(\mathbf{r})$ is a relation-aware modulation function that generates a scaling vector from the relation embedding $\mathbf{r}$.
Finally, the fused multimodal embedding is computed by applying the gate vector to each modality’s hidden projection $\mathbf{h}_{m}$:
@@
where $\mathbf{z}_{m}$ is the $m$-th element gate vector corresponding to modality $m$.
Score Function. After getting the fused multimodal embeddings, we compute triple plausibility scores using the TuckER (Balažević et al., 2019) scoring function:
@@
where $\mathbf{h}_{h}$ and $\mathbf{h}_{t}$ denote the fused embeddings of the head $h$ and tail $t$, $\mathbf{r}_{r}$ denotes the embedding of the relation $r$, and $\times_{n}$ denotes the $n$-mode tensor product.
To train the model, we adopt a binary classification objective that encourages higher scores for positive triples and lower scores for negative ones. Negative samples are generated via uniform negative sampling. The loss function is defined as: